# TreeLearn: A Comprehensive Deep Learning Method for Segmenting Individual Trees from Ground-Based LiDAR Forest Point Clouds

### v2

@@ 7 Conclusion
In this study, we introduced an automatic pipeline for segmenting individual trees from high-resolution forest point clouds, requiring no extensive hyperparameter tuning or data preprocessing in the form of ground and understory removal or terrain normalization. To the best of our knowledge, this work is, along with one concurrent work [37], the first to tackle this task with a deep learning method that operates directly on point cloud data. TreeLearn is based on a 3D U-Net for pointwise feature prediction, followed by clustering and postprocessing to derive instance predictions. As demonstrated in Section 6, our method performs equally well or better than Lidar360 in all aspects, despite being trained on data segmented by it. This highlights the potential of deep learning to replace more complex tree segmentation algorithms to reduce computational demands. Furthermore, our results indicate that major performance increases can be achieved by fine-tuning the network on small amounts of high-quality labeled data. In addition to the segmentation method, we introduced a new benchmark dataset that allows for model training and evaluation of instance segmentation results. This is an important step towards the creation and comparability of powerful deep-learning-based tree segmentation algorithms.
Future research should aim to incorporate TreeLearn in a more general forest inventory framework. Specifically, if adequate training data is available, our method offers the possibility to seamlessly integrate a more fine-grained semantic segmentation into ground, understory, woody parts and leaves, without any changes to the overall pipeline. Furthermore, limitations resulting from the training data used should be addressed. For example, our model was trained with crops with an edge buffer of 11 m where trees with a height of less than 10 m were labeled as non-trees. If trees that are wider than $2*11~{}\text{m}=22~{}\text{m}$ or smaller than 10 m should be correctly recognized, the training data must be adapted accordingly. Furthermore, only beech forests have been used during training. We expect a drop in performance when segmenting point clouds of forests with a highly different structure. Incorporating more diverse data during training or fine-tuning can alleviate this issue. On a positive note, this also means that the model can benefit from any tree segmentation dataset contribution made by the community. In future work, we plan to gather and train on forest data from previous works [46, 10, 47] in an effort to obtain more powerful and generalizable models independent of the characteristics of specific forests or laser scanners. The availability of such models is an important prerequisite for large-scale detailed forest inventory since forest scientists and managers crucially rely on single-tree point clouds for the extraction of single-tree parameters. Such data is urgently needed to plan silviculture activities, to estimate carbon stocks or harvest yields, and for many others tasks.
Appendix A Robustness analysis for grouping radius
@@
The density-based clustering algorithm described in Section 3.4 is controlled by the two hyperparameters $\tau_{\textrm{min}}$ and $\tau_{\textrm{group}}$. The selected minimum number of points for a valid cluster $\tau_{\textrm{min}}=100$ is relatively low. Therefore, there is the risk of incurring false positive tree predictions that do not actually constitute a tree but only a fragment of a tree. To counteract the emergence of false positives, the grouping radius can be increased. However, at some point this might lead to false negative predictions due to tree clusters merging together. To quantify this trade-off between false positives and false negatives, we conducted an experiment where we vary the grouping radius $\tau_{\textrm{group}}$ between 2.5 cm and 70 cm and measure the corresponding detection failures. For the fine-tuned model, the expected tendency for clusters to break apart could be observed for a very small grouping radius of 2.5 cm, while this was not the case for the model trained on noisy labels (Figure 8, green line). This can be attributed to the fact that, in the fine-tuned setting, a single tree-cluster is more likely to be split into distinguishable sub-clusters due to a more precise offset prediction. Furthermore, the expected tendency for clusters to merge could be observed for both settings (Figure 8, red line). However, the number of detection failures is in general very low and not strongly influenced by changes in the grouping radius, which suggests that our method is robust to changes in this hyperparameter. For practical purposes, we recommend setting the grouping radius relatively low (somewhere in the range of 0.1 m to 0.2 m) since false positives can be easily dealt with by simply merging fragments together, while merged trees cannot be that easily recovered.
Figure 8: Number of false positive and false negative tree predictions depending on the grouping radius for (a) the non-fine-tuned model and (b) the fine-tuned model.

### v3

@@ 4 Conclusion
In this study, we introduced an automatic pipeline for segmenting individual trees from high-resolution forest point clouds, requiring no extensive hyperparameter tuning or data preprocessing in the form of ground and understory removal or terrain normalization. TreeLearn is based on a 3D U-Net for pointwise offset prediction, followed by clustering and postprocessing to derive instance predictions. Compared to two concurrent works that also employ offset prediction [41, 42], TreeLearn avoids the need for complicated block-merging heuristics by projecting points towards their respective tree base, and clustering complete instances in a single step on the entire input point cloud. Our method achieves competitive detection and segmentation results compared to recent classical and deep-learning-based tree instance segmentation algorithms. In addition to the segmentation method, we introduced a new benchmark dataset that allows for model training and evaluation of instance segmentation results. Furthermore, we extended two previously published labeled forest point clouds [13, 52] to obtain a consistent input structure for neural network training. This is an important step towards the creation and comparability of powerful deep-learning-based tree segmentation algorithms.
We highlighted several failure cases of our method that arise from difficulties in clearly separating tree instances in the space of offset-shifted coordinates. These should be tackled by future research. Furthermore, we found that using larger amounts of training data steadily increases segmentation performance. This reinforces the importance of making high-quality labeled forest datasets publicly available. A large and diverse data basis is especially important for the development of segmentation models that are applicable to point clouds with diverse laser scanning and forest characteristics. This emerging field of research [42, 71] deserves particular attention. Data-driven segmentation methods like TreeLearn will, in our opinion, play a pivotal role in this endeavour. These methods also offer the possibility to obtain additional output for detailed forest inventory. For example, Xiang et al. [41] demonstrated that a fine-grained semantic segmentation, e.g. into leaves and woody parts, can be directly added to tree instance segmentation models. This functionality can also be seamlessly integrated into TreeLearn without any changes to the overall pipeline. Further potential lies in the integration of tree species classification, which should be investigated by future research. All datasets that were generated in the course of this work as well as the code for training, running and evaluating TreeLearn on arbitrary forest point clouds is made publicly available.
The availability of powerful tree segmentation methods is an important prerequisite for large-scale detailed forest inventory since forest scientists and managers crucially rely on single-tree point clouds for the extraction of single-tree parameters. Such data is urgently needed to plan silviculture activities, to estimate carbon stocks or harvest yields, and for many others tasks.
Code and data availability
The TreeLearn code is available from https://github.com/ecker-lab/TreeLearn. The automatically labeled data, our manually labeled benchmark dataset as well as trained models can be found at https://doi.org/10.25625/VPMPID. The extended versions of the LAUTx and Wytham Woods datasets are available from https://doi.org/10.25625/QUTUWU.
CRediT authorship contribution statement
J. H.: conceptualization, methodology, analysis, writing - original draft, visualization, software; J. v. D.: conceptualization, methodology, analysis, writing - original draft, visualization, software; D. S.: data curation, writing - review and editing; T. K.: writing - review and editing; A. E.: supervision, writing – review and editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT [72] in order to paraphrase text and generate basic code for data processing and figure generation. After using ChatGPT, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix A Datasets
Table 4 contains detailed information on the 19 forest plots from our data. All of the forest plots only contain labeled trees that are at least 10 m high. Therefore, smaller trees were not taken into account in the calculation of the metrics. The tree height was calculated as the difference between the 5th-highest point and the 5th-lowest point of each tree to account for outliers. To obtain the crown diameter, tree points were first projected on the xy-plane. Then a concave hull was fitted on these coordinates. The crown diameter is defined as the farthest distance between two points on this hull. The canopy cover was calculated as the area of the hull. It should be emphasized that all plots except for L1W come with pseudo tree labels generated using the Lidar360 software. The tree parameters were extracted using these labels and should therefore only be treated as a rough approximation of the ground truth. For the L1W dataset, the tree parameters were calculated based on the 156 complete trees that were segmented manually such that they are more accurate. In L1W, the understory is sparse and there are only a few juvenile trees in the area such that the raw forest floor is mostly visible. Further information on the plots is provided by Neudam et al. [56].
Table 4: Detailed information on the 19 forest plots from our data. Tree age, tree height, crown diameter and canopy cover are the mean values across all trees in a plot $\pm$ the standard deviation.
plot name study area $n$ trees plot size ($ha$) tree age (years) tree height ($m$) crown diameter ($m$) canopy cover ($m^{2}$)
A1N Altstedt 466 1.95 141 26.3$\pm$6.9 10.5$\pm$3.6 54.9$\pm$36.5
A1W Altstedt 511 1.42 67 22.0$\pm$5.5 8.8$\pm$2.7 39.5$\pm$23.1
O1N Oppershofen 160 1.22 163 39.2$\pm$9.4 13.4$\pm$4.5 88.3$\pm$53.2
O1W Oppershofen 266 1.46 124 27.5$\pm$10.0 10.7$\pm$4.0 61.5$\pm$38.6
G1N Göttingen 320 1.07 143 30.1$\pm$7.3 9.5$\pm$3.1 45.4$\pm$26.2
G1W Göttingen 450 2.15 124 27.8$\pm$6.5 10.9$\pm$2.7 61.2$\pm$26.6
G2N Göttingen 496 1.56 162 23.9$\pm$7.5 10.0$\pm$4.0 53.5$\pm$39.3
G2W Göttingen 265 1.00 149 22.9$\pm$10.2 9.4$\pm$3.9 49.7$\pm$33.9
G3N Göttingen 355 1.56 137 29.6$\pm$8.4 10.8$\pm$3.3 58.2$\pm$31.0
G3W Göttingen 332 1.48 129 27.1$\pm$8.0 10.8$\pm$3.1 59.4$\pm$29.1
G4N Göttingen 318 1.08 133 27.2$\pm$9.1 10.3$\pm$3.5 54.1$\pm$33.9
G4W Göttingen 585 2.19 130 29.2$\pm$8.6 10.7$\pm$3.5 57.6$\pm$36.2
L1N Lübeck 453 1.69 151 26.3$\pm$8.8 11.1$\pm$3.7 65.2$\pm$41.1
L2N Lübeck 233 1.29 135 30.8$\pm$8.9 12.7$\pm$4.0 81.8$\pm$43.4
L2W Lübeck 401 1.67 92 25.0$\pm$8.3 10.6$\pm$3.4 57.8$\pm$33.9
LG1 Lübeck 407 1.14 128 33.9$\pm$7.6 8.8$\pm$2.7 38.0$\pm$21.5
LG2 Lübeck 332 1.53 127 28.3$\pm$9.0 12.2$\pm$3.9 74.0$\pm$39.0
LG3 Lübeck 315 1.59 149 25.5$\pm$5.7 11.9$\pm$3.4 71.0$\pm$35.8
L1W Lübeck 156 1.16 136 30.0$\pm$6.5 12.6$\pm$3.0 85.2$\pm$36.8
Figure 7: Propagation of segmentation labels to the complete forest point clouds for Wytham Woods (left) and LAUTx (right). Non-tree and unlabeled points were added. Furthermore minor mistakes were corrected as can be seen for the tree bases in LAUTx.
Appendix B Clustering
Figure 8: Clustering of the offset-shifted coordinates on the L1W dataset. (a) shows the initial clustering result that only takes into account points that are close to the tree base and have a sufficiently high verticality (see Section 2.2.4 for details) and (b) shows the results after assigning the remaining points (see Section 2.2.5 for details).
Appendix C Robustness analysis for grouping radius
Figure 9: Number of commission and omission errors depending on the grouping radius for (a) TreeLearn trained with Lidar360 labels and (b) TreeLearn that was additionally fine-tuned.
@@ Appendix C Robustness analysis for grouping radius
The density-based clustering algorithm described in Section 2.2.4 is controlled by the two hyperparameters $\tau_{\textrm{min}}$ and $\tau_{\textrm{group}}$. The selected minimum number of points for a valid cluster $\tau_{\textrm{min}}=100$ is relatively low. Therefore, there is the risk of incurring commission errors if a tree prediction does not actually constitute a tree but only a fragment. To reduce the number of commission errors, the grouping radius can be increased. However, at some point this might lead to omission errors due to tree clusters merging together. To quantify this trade-off between commission and omission errors, we conducted an experiment where we vary the grouping radius $\tau_{\textrm{group}}$ between 2.5 cm and 70 cm and measure the corresponding detection errors on L1W. For the fine-tuned model, the expected tendency for clusters to break apart could be observed for a very small grouping radius of 2.5 cm, while this was not the case for the model trained on noisy labels (Figure 9, orange line). This can be attributed to the fact that, in the fine-tuned setting, a single tree-cluster is more likely to be split into distinguishable sub-clusters due to a more precise offset prediction. Furthermore, the expected tendency for clusters to merge could be observed for both settings (Figure 9, blue line). However, the number of detection failures is in general very low and not strongly influenced by changes in the grouping radius, which suggests that our method is robust to changes in this hyperparameter. For practical purposes, we recommend setting the grouping radius relatively low (somewhere in the range of 0.1 m to 0.2 m) since commission errors can be easily dealt with by simply merging fragments together, while merged trees cannot be that easily recovered. It should be noted that this experiment was conducted with an older version of the model and a slightly different evaluation protocol, which is why the reported error rates differ from what is reported in Section 3.
Appendix D Offset loss
Table 5: Offset loss of TreeLearn models in m. Value in brackets shows the loss when disregarding the z-component.
Test dataset Method Offset Loss ↓
L1W TreeLearn noisy 2.41 (2.27)
TreeLearn small 2.72 (2.30)
TreeLearn mid 2.47 (2.27)
TreeLearn large 2.40 (2.25)
Wytham Woods TreeLearn noisy 6.15 (5.60)
TreeLearn small 6.35 (5.84)
TreeLearn mid 6.20 (5.80)
Table 5 shows the offset loss for the TreeLearn models trained with different amounts of training data. The value of the offset loss when only considering the x- and y-component of the offset vector is also shown since the z-component is ignored during clustering. For the fine-tuned models, a lower offset loss corresponds to a better performance, as is expected. TreeLearn noisy has the lowest offset loss for both L1W and Wytham. However, it should be noted that this model was trained with an older version of our pipeline where the input point cloud was additionally filtered using statistical outlier removal. Therefore, the loss is computed on a different set of points and the better results cannot be clearly interpreted. In general, the offset loss is better on L1W, reflecting the better segmentation performance on this dataset. The validation loss during fine-tuning on plot 16 of the NIBIO dataset was between 0.5 m and 0.7 m for all models. This value is much lower which might be explained by the similarity of the training data to it. Most point clouds used during training have also been captured via UAV. It might also be a consequence of this plot being relatively easy to segment.